consider a group of n= 4 people with the following ages: 16, 18, 20 and 22. Consider samples of size n= 2 from the group. If X is the average age of the two people in a sample, find:
the mean and variance of the sampling distribution of X.
compare these values to the mean and variance of the population.
1
Expert's answer
2021-07-18T08:22:10-0400


We have population values "16, 18, 20, 22," population size "N=4," and sample size "n=2." Thus, the number of possible samples which can be drawn without replacement is



"\dbinom{4}{2}=6"



"mean=\mu=\dfrac{16+18+20+22}{4}=19"

"Variance=\sigma^2"

"=\dfrac{(16-19)^2+(18-19)^2+(20-19)^2+(22-19)^2}{4}=5"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 No & Sample & Mean \\ \hline
 1 & (16, 18) & 17 \\
 \hdashline
 2 & (16, 20) & 18 \\
 \hdashline
3 & (16, 22) & 19 \\
 \hdashline
4 & (18, 20) & 19 \\
 \hdashline
5 & (18, 22) & 20 \\
 \hdashline
6& (20, 22) & 21 \\
 \hdashline
\end{array}"

The sampling distribution of the sample mean "\bar{x}" and its mean and standard deviation are:


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:}
 \bar{x} & f & f(\bar{x})& \bar{x}f(\bar{x})& \bar{x}^2f(\bar{x}) \\ \hline
 17 & 1 & 1/6 & 17/6 & 289/6 \\
 \hdashline
 18 & 1 & 1/6 & 18/6 & 324/6 \\
 \hdashline
 19 & 2 & 2/6 & 38/6 & 722/6 \\
 \hdashline
 20 & 1 & 1/6 & 20/6 & 400/6 \\
 \hdashline
 21 & 1 & 1/6 & 21/6 & 441/6 \\
 \hdashline
 Total & 6 & 1 & 19 & 1088/3 \\
 \hdashline
\end{array}"

"E(\bar{X})=\sum\bar{x}f(\bar{x})=19"

"Var(\bar{X})=\sum\bar{x}^2f(\bar{x})-(\sum\bar{x}f(\bar{x}))^2"

"=\dfrac{1088}{3}-19^2=\dfrac{5}{3}"

"E(\bar{X})=19=\mu"

"Var(\bar{X})=\dfrac{5}{3}=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})=\dfrac{5}{2}(\dfrac{4-2}{4-1})"